# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ 4 Robust Instance Optimality
We first announce our result, which is novel for nonlinear compressed sensing and closely resembles the standard guarantee in linear case (e.g., Proposition 2.1). It is also a formal version of the second informal statement provided in introduction.
@@
Consider the above setting with sufficiently small $\tau_{0}$ and $\zeta_{0}$, and the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ where $\hat{\mathbf{x}}$ is solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8) with
$\varepsilon=C_{1}\tau_{0}+C_{2}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}+C_{3}\sqrt{\frac{s\log({en}/{s})}{m}}$ (4.1)
@@ 4.1 Proof Strategy
To prove Theorem 4.1, we essentially need to run the arguments in Theorems 3.2, 3.3, 3.4 and 3.6 again, along with some additional steps to make certain arguments uniform. To avoid repetition, we will only outline the proof strategies and emphasize the additional techniques.
Without loss of generality, we can assume that the supports of $\bm{\zeta}_{1}$ and $\bm{\zeta}_{2}$ are disjoint. By Remark 3.4 we can consolidate $\bm{\zeta}_{1}$ and $\bm{\zeta}_{2}$ to rewrite $\breve{\mathbf{z}}$ as
@@
where $\bm{\zeta}\in\mathbb{C}^{m}$ satisfies $\|\bm{\zeta}\|_{0}\leq 2\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$. Recall also that we can focus on $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}:=\mathcal{X}$ (see the proof of Theorem 3.2). Further, we let $\tilde{\bm{\tau}}:=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ and write
@@
We then perform the two-step analysis. As we shall see, extra care is needed to ensure that every piece of the prior arguments is uniform for all $\mathbf{x}\in\mathcal{X}$, and the major additional work is on a uniform upper bound on $|\mathcal{J}_{\mathbf{x},\eta_{0}}|=|\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}|$ and a uniform concentration bound on $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$. We proceed in several steps.
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: Corollary 1 yields $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},c_{1})$ for some sufficiently small $c_{1}$, and one can check that the relevant arguments in the proofs of Theorems 3.3 and 3.6 can show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{u}\|_{2}$ and $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{u}\|_{2}$ are sufficiently small. Therefore, it remains to show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{u}\|_{2}$ to be small enough. In the proof of Theorem 3.4, the central idea is the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\tilde{\bm{\tau}}_{1}$ is near-dense and has entries bounded by $\frac{2\tau_{0}}{\eta_{0}}$, and $\tilde{\bm{\tau}}_{2}$ is $(\eta_{0}m)$-sparse with support set $\mathcal{J}_{\mathbf{x},\eta_{0}}$ and has entries bounded by $2$. We obtain the desired result by setting $\eta_{0}$ as a small absolute constant because we can then re-iterate the arguments for small dense noise to treat $\tilde{\bm{\tau}}_{1}$, and these for sparse phase corruption to treat $\tilde{\bm{\tau}}_{2}$. Such idea and most arguments remain valid, while the only notable issue is that the proof of Theorem 3.4 only shows $\|\tilde{\bm{\tau}}_{2}\|_{0}=|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$ for a fixed $\mathbf{x}$ and hence only ensures the existence of the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ for this $\mathbf{x}$. We need to strengthen this step and show that such a decomposition exists for all $\mathbf{x}\in\mathcal{X}$; to that end, we need to show $\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$. This is still true with high probability under $m=\Omega(s\log(\frac{en}{s}))$, as guaranteed by the following lemma. See a more refined statement and the proof in Appendix B.2.
@@
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: By (4.3) and $\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}=\mathbf{e}_{1}$, we seek to bound
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$ $\leq\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}^{\star}\|_{2}+\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}^{\star}\|_{2}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}^{\star}\|_{2}$
$=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big{(}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}\|_{2}+\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}+\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\Big{)}.$
Recall that we repeatedly use the concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$; e.g., see (3.7), (3.13), (3.21)). While this follows from sub-Gaussian concentration for a fixed $\mathbf{x}$, we need to strengthen it to uniform concentration over all $\mathbf{x}\in\mathcal{X}$. The following lemma shows that $\sup_{\mathbf{x}\in\mathcal{X}}|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough under $m=\Omega(s\log(\frac{en}{s}))$. Up to some simple modifications (from $\mathbb{R}$ to $\mathbb{C}$), the proof is identical to that of [42, Lem. 2.1], and is hence omitted.
Lemma 4.2.
Now, for $\mathbf{w}=\bm{\tau}_{(2)},\tilde{\bm{\tau}},\bm{\zeta}$, it remains to bound
$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\mathbf{w}}\mathbf{x}\|_{2}\leq\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}.$
We have
$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\tau}_{(2)}}\mathbf{x}\|_{2}\leq\tau_{0}\left[\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\right]=O(\tau_{0})$
due to (2.10). For $\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$, we also have $\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}=O(\tau_{0})$ as our original arguments in (3.14)–(3.16) are already uniform over $\mathbf{x}\in\mathcal{X}$.
The bound on $\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}$ is a bit more tricky. From (3.22) we have
$\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq C_{1}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2},$
and for a fixed $\mathbf{x}$ we achieve the bound $\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}=O(\sqrt{\zeta_{0}\log(1/\zeta_{0})})$ through (3.23). In contrast, here we need a uniform bound over $\mathbf{x}\in\mathcal{X}$, and we use Lemma 2.1 to obtain
$\sup_{\mathbf{x}\in\mathcal{X}}\|\mathbf{A}_{\bm{\zeta}}\mathbf{x}\|_{2}\leq C_{1}\sup_{\mathbf{x}\in\mathcal{X}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=\zeta_{0}m\end{subarray}}\left(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\right)^{1/2}=O\left(\sqrt{\zeta_{0}\log(e/\zeta_{0})}+\sqrt{\frac{s\log(en/s)}{m}}\right).$
Therefore, a new term $O(\sqrt{\frac{s}{m}\log(\frac{en}{s})})$ arises in the error horizon, and consequently it also appears in our choice of the noise level (4.1).

### v3

@@ 4 Robust Instance Optimality
We first announce our result, which is novel for nonlinear compressed sensing and closely resembles the standard guarantee in linear case (e.g., Proposition 2.1). It is also the formal version of the second informal statement provided in introduction.
@@
Consider the above setting with $\max\{\tau_{0},\zeta_{0}\}\leq c_{0}$ for some sufficiently small $c_{0}$, and the estimator $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ where $\hat{\mathbf{x}}$ is solved from $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1};\varepsilon)$ in (1.8) with
$\varepsilon=C_{1}\tau_{0}+C_{2}\sqrt{\zeta_{0}\log({e}/{\zeta_{0}})}+C_{3}\sqrt{\frac{s\log({en}/{s})}{m}}\mathbbm{1}(\zeta_{0}>0)$ (4.1)
@@ 4.1 Proof Strategy
Without loss of generality, we can assume that the supports of $\bm{\zeta}_{(1)}$ and $\bm{\zeta}_{(2)}$ are disjoint. By Remark 3.5 we can consolidate $\bm{\zeta}_{(1)}$ and $\bm{\zeta}_{(2)}$ to rewrite $\breve{\mathbf{z}}$ as
@@
where $\bm{\zeta}\in\mathbb{C}^{m}$ satisfies $\|\bm{\zeta}\|_{0}\leq 2\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$. Further, we let $\tilde{\bm{\tau}}:=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ and write
@@
Recall also that we can focus on $\mathbf{x}\in\mathbb{B}_{1}^{n}(\sqrt{2s})\cap\mathbb{S}^{n-1}:=\mathcal{X}$ (see the proof of Theorem 3.2).
We now make the following claim.
Claim 1.
To prove Theorem 4.1, all that remains is to establish
[…]
$\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{9},$ (4.5)
$\frac{13}{14}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}\leq\frac{15}{14}\,,\quad\forall\mathbf{x}\in\mathcal{X}.$ (4.6)
Proof.
To see this claim, we note that (4.4)–(4.5) strengthen the conditions in Lemma 3.2 by making them uniform over $\mathbf{x}\in\mathcal{X}$. Since the RIP of $\mathbf{A}_{\mathbf{z}}$ in Corollary 1 used in Equation (3.9) is uniform over all $\mathbf{x}\in\mathcal{X}$ with the promised probability, the argument of Lemma 3.2 applies to all $\mathbf{x}\in\mathcal{X}$ and yields
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{10\sigma_{\ell_{1}}(\mathbf{x};\Sigma^{n}_{s})}{\sqrt{s}}+\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}\,,\quad\forall\mathbf{x}\in\mathcal{X}.$
(Note that the term $\frac{10\sigma_{\ell_{1}}(\mathbf{x};\Sigma^{n}_{s})}{\sqrt{s}}$ appears since $\mathbf{x}$ may not be in $\Sigma^{n,*}_{s}$.) Combining with (4.6) yields (4.2). ∎
All that remains is to show (4.4)–(4.6) separately. To this end, the arguments in Theorems 3.2, 3.3, 3.4, and 3.6 must be revisited, together with additional techniques to ensure uniformity over all $\mathbf{x}\in\mathcal{X}$. To avoid repetition, we only outline the proof strategy and highlight the additional techniques.
Show (4.6). Previously, we use (3.2) to achieve this for a fixed $\mathbf{x}$. Here, we need to strengthen it to uniform concentration over all $\mathbf{x}\in\mathcal{X}$. The following lemma shows that $\sup_{\mathbf{x}\in\mathcal{X}}|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ is small enough under $m=\Omega(s\log(\frac{en}{s}))$. Up to some simple modifications (from $\mathbb{R}$ to $\mathbb{C}$), the proof is identical to that of [44, Lem. 2.1], and is hence omitted.
Lemma 4.1.
Show (4.5). By (4.3) and triangle inequality, it suffices to show, for $\mathbf{w}=\bm{\tau}_{(2)},\tilde{\bm{\tau}},\bm{\zeta}$, that
$\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{27}.$ (4.7)
For $\mathbf{w}=\bm{\tau}_{(2)}$ satisfying $\|\bm{\tau}_{(2)}\|_{\infty}\leq\tau_{0}$, re-iterating the arguments in Equations (3.10)–(3.12) shows that the left-hand side of (4.7) is bounded by $O(\tau_{0})$, and thus smaller than $\frac{1}{27}$ under small enough $\tau_{0}$. For $\mathbf{w}=\bm{\zeta}$ obeying $\|\bm{\zeta}\|_{0}\leq 2\zeta_{0}m$ and $\|\bm{\zeta}\|_{\infty}\leq 2$, the arguments in Equations (3.18)–(3.20) are sufficient to show that the left-hand side of (4.7) is bounded by $O(\sqrt{\frac{s}{m}\log(\frac{en}{s})})+\sqrt{\zeta_{0}\log(1/\zeta_{0})})$, and therefore is smaller than $\frac{1}{27}$ under small enough $\zeta_{0}$ and $m\gtrsim s\log(\frac{en}{s})$.
It is more tricky to establish (4.7) for $\mathbf{w}=\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$ for $\bm{\tau}_{(1)}\leq\tau_{0}$. To this end, we first revisit the argument in Theorem 3.4: Define $\eta_{0}=C^{*}c_{0}$ for some suitably large absolute constant $C^{*}$ and the index set $\mathcal{J}_{\mathbf{x},\eta_{0}}:=\{i\in[m]:|\bm{\Phi}_{i}^{*}\mathbf{x}|\leq\eta_{0}\}$, the key ingredient is a decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\tilde{\bm{\tau}}_{1}$ and $\tilde{\bm{\tau}}_{2}$ accommodate the entries in $\mathcal{J}_{\mathbf{x},\eta_{0}}^{c}$ and $\mathcal{J}_{\mathbf{x},\eta_{0}}$, respectively; note that the entries of $\tilde{\bm{\tau}}_{1}$ is bounded by $\frac{2\tau_{0}}{\eta_{0}}\leq\frac{2c_{0}}{C^{*}c_{0}}=\frac{2}{C^{*}}$, which is suitably small under a large $C^{*}$, and thus $\tilde{\bm{\tau}}_{1}$ can be treated similarly to $\bm{\tau}_{(2)}$ by the arguments in Equations (3.10)–(3.12); under the high-probability event $\{|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\}$ (due to Chernoff bound), $\tilde{\bm{\tau}}_{2}$ satisfies $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$ and $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$; since $\eta_{0}=C^{*}c_{0}$ is small enough (under absolute constant $C^{*}$ and small enough $c_{0}$), $\tilde{\bm{\tau}}_{2}$ can be treated similarly to $\bm{\zeta}$ via the arguments in Equations (3.18)–(3.20). The proof strategy remains valid, with the only gap being that the event $\{|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\}$ for a fixed $\mathbf{x}$ needs to be strengthened to $\{\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\}$, in order to guarantee that the decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ for some $(\eta_{0}m)$-sparse $\tilde{\bm{\tau}}_{2}$ exists for all $\mathbf{x}\in\mathcal{X}$. The following lemma shows that $\sup_{\mathbf{x}\in\mathcal{X}}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m$ holds with high probability under $m=\Omega(s\log(\frac{en}{s}))$. See a more refined statement and the proof in Appendix B.2.
@@
Show (4.4). By (4.3), triangle inequality, $\sup_{\mathbf{x}\in\mathcal{X}}\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\leq\frac{14}{13}$ from (4.6) and $\varepsilon$ in (4.1), it suffices to show
$\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\mathbf{w}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\mathbf{w}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}$
$=O\left(\tau_{0}+\sqrt{\zeta_{0}\log(e/\zeta_{0})}+\sqrt{\frac{s\log(en/s)}{m}}\mathbbm{1}(\zeta_{0}>0)\right)$ (4.8)
For $\mathbf{w}=\bm{\tau}_{(2)}$, $\tilde{\bm{\tau}}$ and $\bm{\zeta}$. For $\mathbf{w}=\bm{\tau}_{(2)}$ obeying $\|\bm{\tau}_{(2)}\|_{\infty}\leq\tau_{0}$, we use (2.10) and (4.6) to achieve
$\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\bm{\tau}_{(2)}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}_{(2)}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\leq\tau_{0}\left(\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\right)=O(\tau_{0}).$
For $\mathbf{w}=\tilde{\bm{\tau}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau}_{(1)})-\operatorname{sign}(\bm{\Phi}\mathbf{x})$, our original calculations in Equations (3.22)–(3.24) suffice to show that the left-hand side of (4.8) is bounded by $O(\tau_{0})$. When showing (4.8) with $\mathbf{w}=\bm{\zeta}$, however, an additional term $\sqrt{\frac{s}{m}\log(\frac{en}{s})}\mathbbm{1}(\zeta_{0}>0)$ arises compared to the non-uniform analysis in Theorem 3.6: For a fixed $\mathbf{x}$, Equations (3.29)–(3.30) establish
$\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\leq C_{1}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=2\zeta_{0}m\end{subarray}}\Big(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\Big)^{1/2}\leq C_{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}\,;$
in contrast, here we establish a uniform bound on $\mathbf{x}\in\mathcal{X}$ by using Lemma 2.1 to obtain
$\sup_{\mathbf{x}\in\mathcal{X}}\frac{|\Re(\bm{\zeta}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\sup_{\mathbf{x}\in\mathcal{X}}\frac{\|\Im(\operatorname{diag}(\bm{\zeta}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}$
$\leq C_{1}\sup_{\mathbf{x}\in\mathcal{X}}\max_{\begin{subarray}{c}I\subset[m]\\ |I|=2\zeta_{0}m\end{subarray}}\Big(\frac{1}{m}\sum_{i\in I}|\bm{\Phi}_{i}^{*}\mathbf{x}|^{2}\Big)^{1/2}\leq C_{2}\sqrt{\zeta_{0}\log(e/\zeta_{0})}+C_{3}\sqrt{\frac{s\log(en/s)}{m}}\mathbbm{1}(\zeta_{0}>0),$
yielding (4.8) with $\mathbf{w}=\bm{\zeta}$.